OpenStudy (anonymous):

Graph the polar curve.. r^2 = sin2(o).. how do I simplify this?

OpenStudy (tkhunny):

Is that \(\sin^{2}(\theta)\;or\;\sin(2\theta)\;or\;something\;else?\)

OpenStudy (anonymous):

\[r^2=\sin2 \theta \]

OpenStudy (tkhunny):

What's stopping you from picking various values of \(\theta\)?

OpenStudy (anonymous):

The \[r^2\]

OpenStudy (tkhunny):

\(\theta = 0 \implies \sin(2\theta) = \sin(2(0)) = \sin(0) = 0\implies r = 0\) Okay, no you do \(\theta = \pi/12\).

OpenStudy (anonymous):

theta = pi/12 sin(2*pi/12)= 1/2 do I square root 1/2 or..?

OpenStudy (tkhunny):

Why would you do anything else? \(r^{2} = 1/2 \implies r = 1/\sqrt{2}\;or\;r = -1/\sqrt{2}\) Let's roll on. Try \(\theta = \pi/8\).